### v2

# Joonkyo Hong    Davide Luparello

Abstract
We assess the replicability of Orr (2022)’s method for estimating within-plant productivity across product lines, which combines demand estimation with cost minimization. The original study uses input price shocks in other output markets as instrumental variables, with exclusion restrictions based on downstream purchase shares. Using the original dataset of Indian machinery producers from 2000–2007, we successfully reproduce the main productivity patterns and demonstrate their robustness to variations in the exclusion threshold. However, extending the analysis to 2010–2019 and 2000–2019 periods yields inadmissible demand systems. Nevertheless, the main results remain robust when calibrating demand parameters to Orr (2022)’s original estimates.
Keywords: Replication, Robustness, Productivity
JEL codes: D22, D24, L64
1 Introduction
Productivity variation across product lines is a critical determinant of overall efficiency (Eckel and Neary, 2010; Bernard et al., 2010; Mayer et al., 2021), yet empirical progress has been hampered by the unavailability of producer-product-level input data. Orr (2022) addresses this challenge by developing a two-stage estimation procedure: the first stage recovers product-specific input allocations by combining the estimation of product demand systems and cost minimization conditions, while the second stage estimates the product-level production function. Analyzing data on Indian machinery producers from 2000–2007, Orr (2022) establishes two main findings: eliminating the lowest-TFPR product increases plant-level TFPR by 10–65 percent, and a one standard deviation decrease in product-level TFPR raises the probability of product discontinuation by 6 percentage points.
Orr (2022)’s methodology confronts a well-known endogeneity concern: product prices may correlate with demand through unobservable characteristics such as quality and appeal. To address it, Orr (2022) constructs instrumental variables using average input price growth facing plants in non-machinery industries, under the identifying assumption that these input price shocks remain orthogonal to machinery industry demand. This assumption fails when machinery producers constitute a substantial share of downstream demand for a given input. To preserve instrument validity, Orr (2022) excludes input codes from single-product producers when machinery-related plants account for more than 30 percent of purchases of that specific input code.
In this paper, we first reproduce the original results. Since the raw administrative data were not included in the replication package, we obtained them independently from the Indian Ministry of Statistics and Programme Implementation (MOSPI). Our estimates closely match the published results, with only minor numerical discrepancies that do not affect the study’s main conclusions. Next, we conduct a sensitivity analysis of the IV construction threshold (30 percent). We find that instruments strength and estimated demand parameters are sensitive to threshold values. Nonetheless, the main results remain remarkably stable. Finally, we extend the analysis to later years (2010–2019) and the full sample (2000–2019). Across all IV thresholds, the proposed instruments yield inadmissible demand systems. We therefore calibrate the demand parameters to Orr (2022)’s original estimates to recover within-plant productivity measures. The main findings remain robust.
The paper proceeds as follows. Section 2 outlines the main methodological innovation in Orr (2022). Section 3 presents our computational reproduction. Section 4 investigates the sensitivity of demand estimates to IV construction thresholds. Section 5 extends the analysis to later periods. Section 6 concludes.
2 Central Idea: The Use of First Order Conditions
In the literature, unit-level productivity is often recovered as production function estimation residuals. To estimate plant-product level productivity, Orr (2022) employs the following production function [^1] Specified in equation (18) of his original paper.:
$Y_{it}^{j}=\exp(\omega_{it}^{j})(L_{it}^{j})^{\beta_{L}}(K_{it}^{j})^{\beta_{K}}(M_{it}^{j})^{\beta_{M}},$ (1)
where $Y_{it}^{j}$ is the total output of product $j$ by plant $i$ in period $t$; $L_{it}^{j}$, $K_{it}^{j}$, and $M_{it}^{j}$ are the respective labor, capital, and material inputs used to produce product $j$; and $\omega_{it}^{j}$ represents total factor productivity for producing product $j$ (TFPQ).
This production function approach requires input allocations across product lines $(L_{it}^{j},K_{it}^{j},M_{it}^{j})$, yet datasets typically provide only plant-level inputs $(L_{it},K_{it},M_{it})$. Orr (2022) resolves this limitation by exploiting the first-order conditions of cost minimization to infer unobserved input shares across products. This approach yields the following fundamental relationship [^2] Formalized in equation (7) of the original paper.:
$S_{it}^{j}=\frac{MC_{it}^{j}Y_{it}^{j}}{\sum_{j\in\mathbb{Y}_{it}}MC_{it}^{j}Y_{it}^{j}},$ (2)
where $S_{it}^{j}$ denotes the input share for product $j$, $Y_{it}^{j}$ represents the output of product $j$, $MC_{it}^{j}$ captures the marginal cost of producing $j$, and $\mathbb{Y}_{it}$ defines the set of products produced.
With $Y_{it}^{j}$ observable in the data, equation (2) enables product-line estimation of the production function, permitting recovery of plant-product productivity conditional on knowledge of $MC_{it}^{j}$. Calculating $MC_{it}^{j}$ requires first estimating the output demand system using plant-product output and prices. When the resulting price elasticity matrix proves invertible, researchers derive $MC_{it}^{j}$ under explicit market conduct assumptions, such as static Bertrand-Nash competition.
In his empirical application, Orr (2022) employs a nested logit demand system [^3] As shown in equation (20) of the original paper.:
$rs_{it}^{j}-rs_{t}^{0}=(1-\sigma)rs_{it}^{j|g(j)}-\alpha p_{it}^{j}+\eta_{it}^{j},$ (3)
where $rs_{it}^{j}=\ln{\frac{R_{it}^{j}}{I_{t}^{h(j)}}}$, with $R_{it}^{j}$ denoting revenue from product $j$ of plant $i$ at time $t$ and $I_{t}^{h(j)}$ capturing total revenue of the 3-digit ASICC code $h(j)$. The term $rs_{it}^{j|g(j)}=\ln{\frac{R_{it}^{j}}{\Lambda_{t}^{g(j)}}}$, where $\Lambda_{t}^{g(j)}$ denotes total revenue of the 5-digit ASICC code $g(j)$. The variable $p_{it}^{j}$ captures the logged output price for product $j$, while $\eta_{it}^{j}$ reflects product appeal. The price elasticity matrix proves invertible if and only if $\alpha>0$ and $0<\sigma<1$.
3 Computational Reproducibility
Using the replication package from Orr (2022)444 Our replication process identifies an issue in the provided package, specifically in asicc_code_cleaning.do, where line 130 excludes plants in industries 74 and 75 due to missing three-digit codes in the MOSPI ASICC09 file. We revise the do file to address this discrepancy and successfully reproduce all tables and figures from Orr (2022). We thank Scott Orr for clarifying the ASICC09 file structure and providing a modified version compatible with the original code. , we reproduce all main results [^5] Results of this reproduction exercise appear in Online Appendix, Section A.. The package provides data cleaning procedures while excluding raw datasets and analytical files. We obtain raw data for Indian manufacturing plants in the machinery sector (2000–2020) from the Ministry of Statistics and Programme Implementation and focus on the 2000–2007 sample, following the original study’s methodology.
Our reproduction exercise successfully replicates the demand estimates from Orr (2022) while revealing minor discrepancies in production function coefficients, detailed in Table 1. GMM estimation generates output elasticities of 0.617 for labor, 0.239 for capital, and 0.223 for materials, contrasting with the original preferred estimates of 0.626, 0.236, and 0.217 from column (3) [^6] These deviations likely originate from computational differences in STATA’s Mata matrix inversion functions across software versions and operating systems, which influence marginal cost calculations and subsequent input allocation decisions.. These minor variations preserve the main findings both qualitatively and quantitatively.
Table 1: Computational Reproduction of Columns 2, 3, and 4 in Table 4 of Orr (2022): Cobb-Douglas Production Function Estimates
GMM
(1) (2) (3)
Original Rep. Original Rep. Original Rep.
Study Study Study
$\beta_{L}$ 0.331 0.325 0.321 0.315 0.626 0.617
(0.192) (0.192) (0.191) (0.191) (0.261) (0.261)
$\beta_{K}$ 0.101 0.106 0.097 0.102 0.236 0.239
(0.082) (0.082) (0.081) (0.081) (0.099) (0.099)
$\beta_{M}$ 0.790 0.789 0.806 0.806 0.217 0.223
(0.191) (0.191) (0.186) (0.186) (0.352) (0.352)
Ins.
$(Z_{t}^{g(j)},Z_{t-1}^{g(j)})$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
$m_{it-1}$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Observations 3,620 3,620 3,620 3,620 3,620 3,620
  • Notes: Columns 1, 2, and 3 present computational reproductions (Rep.) of Columns 2, 3, and 4, respectively, from Table 4 of Orr (2022) (Original Study). Calculations employ the 2000–2007 Indian ASI dataset provided by the Ministry of Statistics and Programme Implementation (MOSPI). Each observation represents a plant-product combination at the 5-digit ASICC variety level. The sample restricts analysis to producers within the machinery, equipment, and parts industry. The dependent variable captures the log quantity of product $j$ produced by plant $i$. Parentheses report plant-level block bootstrapped standard errors with 1,000 replications.
4 Robustness of Threshold for Instrument Construction
Orr (2022) excludes input codes utilized by single-product producers when machinery, equipment, and parts producers account for more than 30% of revenue derived from purchases of these input codes. This exclusion criterion preserves instrument validity. As Orr (2022) argues, “input price variation should be driven by demand and supply shocks in other industries that are orthogonal to machinery demand shocks or general changes in machinery quality by product code. This requires that changes in average input prices not be driven by machinery demand.”
The IV construction threshold substantially affects demand estimates. We replicate the analysis across thresholds ranging from 0 to 100% in 1% increments and present results in Figure 1. First, the threshold directly determines sample size: lower thresholds reduce observations while higher thresholds retain more data but increase endogeneity risk. Panel 1(d) displays this variation in sample size.
Second, Panels 1(a) and 1(b) present the price and revenue share parameter estimates. Excluding boundary cases near 0% and 100%, the price coefficient consistently yields a downward-sloping demand curve, while the revenue share coefficient produces nest parameter values within the theoretical unit interval. Significance varies markedly across thresholds, achieving 90% confidence only within the 20%-40% range.
Finally, Panel 1(c) documents instrument strength using Sanderson-Windmeijer F-statistics. Price instruments attain optimal performance within the 20%-40% threshold range, yielding F-statistics between 10 and 17. Outside this interval, F-statistics deteriorate to 1-6. Revenue share instruments demonstrate higher stability, sustaining F-statistics of 8-11 throughout the 20%-40% threshold range. Online Appendix B, Figure A3 presents additional first-stage coefficient estimates.
Given our estimation of downward-sloping demand curves across nearly the entire threshold interval, we examine result sensitivity regarding plant-level efficiency gains from marginal variety elimination and discontinuation probabilities following product-level productivity declines. Figure 2 demonstrates that these findings exhibit remarkable robustness across threshold values.
Under the nonparametric specification from Orr (2022), Panel 2(a) documents that plants discontinuing their worst-performing products realize plant-level revenue efficiency gains spanning 10–15% at lower bounds and reaching 60–80% at upper bounds, with peaks approaching 100% for thresholds of 60–70%. Panel 2(b) confirms that one standard deviation declines in product-level revenue-based total factor productivity increase discontinuation probabilities by 6–7 percentage points across all threshold values. [^7] More detailed results are reported in Figures A1-A2 and Tables A3 in Online Appendix B.
5 Replication Using Different Sample Periods
We apply the methodology of Orr (2022) to (1) the ASI sample for 2010–2019 and (2) the full sample spanning 2000–2019888For comparability, we convert post-2010 product codes (NPCMS) to ASICC09 using the official ASICC09–NPCMS concordance from MOSPI. We report summary statistics for the two samples in Online Appendix C.. We begin by evaluating the performance of the proposed IVs but consistently fail to obtain admissible estimates across the full range of threshold values. Figures 3 and 4 present the demand estimates for the 2010–2019 and 2000–2019 samples, respectively, along with first-stage F-statistics and the number of observations at each threshold. In both samples, the estimated price and nest parameters lie outside the invertible regions across all thresholds. Moreover, our estimates yield no statistically significant results. Consistent with these findings, the instruments exhibit weak power throughout, with first-stage F-statistics never exceeding the conventional threshold of 10.
Although demand estimates from different periods violate the invertibility condition, we assess the robustness of Orr (2022)’s main findings by recovering within-firm input allocation shares using the original estimates and estimating the production function [^9] Production function parameter estimates appear in Online Appendix Section C, Tables A5 and A6.. We then quantify efficiency gains from eliminating the least productive products and evaluate the marginal effect of a one standard deviation decline in revenue-based productivity on product discontinuation probability.
Table 2 presents the main results from this robustness exercise. The patterns exhibit remarkable consistency across sample periods. Removing the lowest-performing products generates substantial increases in plant-level revenue-based productivity: 8.82%–61.60% for the original sample, 1.13%–67.13% for 2011–2020, and 9.94%–65.59% for 2001–2020. The marginal effects of a one standard deviation decline in revenue-based productivity on the probability of product discontinuation display similar stability across periods, with magnitudes of 6.65, 9.22, and 7.96 percentage points, respectively [^10] For more detailed results, please refer to Figures A7-A10 and Tables A7 and A8 in Online Appendix C..
Table 2: Baseline Within-Firm Productivity Dispersion Implications
Period 2000-2007 2010-2019 2000-2019
Plant-level Efficiency Growths (%) [8.82, 61.60] [1.13, 67.13] [9.94, 65.59]
Marginal Effect of 1 SD TFPR Decline 6.65 9.22 7.96
on Product-Dropping (pp) (2.20) (1.53) (2.08)
  • Notes: This table documents robustness checks on the baseline within-firm productivity dispersion findings from Orr (2022). The first column presents lower and upper bounds of plant-level efficiency gains from eliminating marginal varieties of lowest-performing products, alongside marginal effects of one standard deviation declines in revenue-based productivity on product exit probabilities, employing 2000–2007 ASI data from India’s Ministry of Statistics and Programme Implementation (MOSPI). Columns two and three replicate this analysis using 2010–2019 and 2000–2019 ASI samples, respectively. Parentheses report plant-level semi-parametric block bootstrap standard errors based on the parametric distribution of estimated demand parameters from the original sample (1,000 replications).
6 Conclusion
This paper evaluates the reproducibility and replicability of the methodology proposed by Orr (2022) for estimating within-plant productivity across product lines. We reproduce the original results and document that demand estimates are highly sensitive to the thresholds used in constructing instrumental variables. When we extend the analysis to alternative sample periods (2010–2019 and 2000–2019), the demand estimation procedure becomes unstable, requiring calibration of key parameters from the original sample. Nonetheless, the study’s main findings prove robust: the estimated efficiency gains from discontinuing underperforming products, as well as the relationship between productivity declines and product exit probabilities, remain consistent across sample periods.
Appendix A Computational Reproducibility: Additional Results
Table A1: Demand Estimates
Estimates: OLS IV
$p^{j}_{it}$ 0.007 -0.220
(0.002) (0.116)
$rs^{j|g(j)}_{it}$ 0.946 0.621
(0.004) (0.306)
Observations 64,917 64,917
First Stage: $p^{j}_{it}$ $rs^{j|g(j)}_{it}$
$Z^{g(j)}_{t}$ 0.284 0.136
(0.097) (0.041)
$Z^{-jg}_{it}$ 0.322 -0.186
(0.198) (0.108)
Observations 64,917 64,917
  • Notes: This table replicates Table 3 from Orr (2022). The calculations employ the 2000–2007 Indian Annual Survey of Industries (ASI) dataset from the Ministry of Statistics and Programme Implementation (MOSPI). All regressions control for firm age and include indicator variables for product code, year, state, census status, rural locations, organization type, ownership structure, and number of products. The Sanderson-Windmeijer F-statistics equal 16.36 for $p^{j}_{it}$ and 10.07 for $rs^{j|g(j)}_{it}$. Standard errors cluster two-way by plant and product code.
Table A2: Computational Reproduction of Columns 2, 3, and 4 in Table 4 of Orr (2022): Cobb-Douglas Production Function Estimates
GMM
(1) (2) (3)
Original Rep. Original Rep. Original Rep.
Study Study Study
$\beta_{L}$ 0.331 0.325 0.321 0.315 0.626 0.617
(0.192) (0.192) (0.191) (0.191) (0.261) (0.261)
$\beta_{K}$ 0.101 0.106 0.097 0.102 0.236 0.239
(0.082) (0.082) (0.081) (0.081) (0.099) (0.099)
$\beta_{M}$ 0.790 0.789 0.806 0.806 0.217 0.223
(0.191) (0.191) (0.186) (0.186) (0.352) (0.352)
$\rho^{74}$ 0.757 0.757 0.747 0.747 0.842 0.841
(0.222) (0.222) (0.197) (0.197) (0.199) (0.186)
$\rho^{75}$ 0.657 0.658 0.661 0.661 0.670 0.670
(0.082) (0.082) (0.078) (0.078) (0.068) (0.068)
$\rho^{76}$ 0.651 0.652 0.653 0.653 0.623 0.623
(0.098) (0.098) (0.104) (0.104) (0.079) (0.080)
$\rho^{77}$ 0.420 0.420 0.422 0.422 0.541 0.540
(0.062) (0.062) (0.060) (0.060) (0.087) (0.088)
$\rho^{78}$ 0.194 0.195 0.181 0.182 0.569 0.568
(0.319) (0.318) (0.265) (0.264) (0.651) (0.658)
RTS 1.222 1.220 1.224 1.223 1.078 1.078
(0.084) (0.084) (0.080) (0.080) (0.113) (0.112)
Ins.
$(Z_{t}^{g(j)},Z_{t-1}^{g(j)})$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
$m_{it-1}$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Observations 3,620 3,620 3,620 3,620 3,620 3,620
  • Notes: Columns 1, 2, and 3 present computational reproductions (Rep.) of Columns 2, 3, and 4, respectively, from Table 4 of Orr (2022) (Original Study). Calculations employ the 2000–2007 Indian ASI dataset provided by the Ministry of Statistics and Programme Implementation (MOSPI). Each observation represents a plant-product combination at the 5-digit ASICC variety level. The sample restricts analysis to producers within the machinery, equipment, and parts industry. The dependent variable captures the log quantity of product $j$ produced by plant $i$. Parentheses report plant-level block bootstrapped standard errors with 1,000 replications.
Figure A1: Negative Correlation between TFPQ and Product Appeal
  • Notes: This figure replicates Figure 1 from Orr (2022). Appeal refers to $\eta^{j}_{it}$, while TFPQ refers to $\omega^{j}_{it}$. The calculations employ the 2000–2007 Indian Annual Survey of Industries (ASI) dataset from the Ministry of Statistics and Programme Implementation (MOSPI). Both variables exhibit standardization to achieve zero mean and unit variance within each product code. Panel (1) encompasses all products while Panel (2) restricts analysis to single-product plants. Plant-level block bootstrapped standard errors appear in parentheses (1,000 replications).
Figure A2: Plant-Level Revenue Efficiency Growth After Removing Marginal Varieties
  • Notes: This figure replicates Figure 2 from Orr (2022). The figure displays the implied log revenue efficiency growth from removing a plant’s lowest efficiency variety according to core-competence technology. The calculations employ the 2000–2007 Indian Annual Survey of Industries (ASI) dataset from the Ministry of Statistics and Programme Implementation (MOSPI). The y-axis denotes the change in plant-level log revenue efficiency after dropping a plant’s worst product.
Table A3: Product-Dropping Probits (3,472 Observations)
(1) (2) (3) (4) (5) (6) (7) (8)
$E^{j}_{it}$ -0.0666
$\bar{P}^{g}_{t}$ -0.0318 -0.0253
(0.0266) (0.0271)
$\bar{A}^{j}_{it}$ -0.0403 -0.0390
(0.0195) (0.0199)
$\bar{C}^{j}_{it}$ 0.0027 -0.0080
(0.0173) (0.0209)
$\hat{P}^{g}_{t}$ -0.0573 -0.0894
(0.0218) (0.0245)
$\hat{A}^{j}_{it}$ -0.0797 -0.1278
(0.0194) (0.0278)
$\hat{C}^{j}_{it}$ 0.0087 -0.0700
(0.0177) (0.0370)
  • Notes: This table replicates Table 9 from Orr (2022), presenting marginal effects from a probit regression where the dependent variable indicates whether a plant discontinues production of a specific product in the subsequent period. The analysis uses plant–product observations from India’s Annual Survey of Industries (ASI) dataset spanning 2000–2007, obtained from the Ministry of Statistics and Programme Implementation (MOSPI). The sample includes all producers within the machinery, equipment, and parts industry, with each observation representing a unique plant–product combination at the 5-digit ASICC classification level. Standard errors (in parentheses) are clustered at the plant level using block bootstrap with 1,001 replications.
Appendix B Robustness of Threshold for Instrument Construction: Additional Results
Figure A3: Sensitivity of First Stage Demand Estimates to Instrument Construction Threshold
(a) First Stage Coefficient of $p^{j}_{it}$ on $Z^{g(j)}_{t}$
(b) First Stage Coefficient of $rs^{j|g(j)}_{it}$ on $Z^{g(j)}_{t}$
(c) First Stage Coefficient of $p^{j}_{it}$ on $Z^{-jg}_{it}$
(d) First Stage Coefficient of $rs^{j|g(j)}_{it}$ on $Z^{-jg}_{it}$
Notes: This figure displays changes in first-stage estimates reported in Table 3 of Orr (2022) as the instrument construction threshold varies in 0.01 increments. The y-axis presents distinct measures across panels: Panel (a) shows the estimated first-stage coefficient of $p^{j}_{it}$ on $Z^{g(j)}_{t}$, Panel (b) displays the estimated first-stage coefficient of $rs^{j|g(j)}_{it}$ on $Z^{g(j)}_{t}$, Panel (c) reports the estimated first-stage coefficient of $p^{j}_{it}$ on $Z^{-jg}_{it}$, and Panel (d) reports the estimated first-stage coefficient of $rs^{j|g(j)}_{it}$ on $Z^{-jg}_{it}$. All panels include 90% confidence intervals constructed using robust standard errors clustered by plant and product. A vertical dashed line marks the threshold employed in Orr (2022) (0.3). The calculations utilize the 2000–2007 Indian Annual Survey of Industries (ASI) dataset from the Ministry of Statistics and Programme Implementation (MOSPI).
Appendix C Robust Replication Using Different Sample Periods: Additional Results
Table A4: Plant-Year Summary Statistics: Machinery, Equipment, and Parts
Panel A: 2000–2007 Original Sample (20,701 Observations) Variable Mean SD Min Max Median Log Labor 8.92 1.44 3.26 14.54 8.65 Log Capital Stock 15.15 2.25 -.09 23.48 14.94 Log Materials 6.91 3.29 -3.91 22.99 6.39 No. of varieties 1.55 1.31 1 28 1 Multiproduct .26 .44 0 1 0
Panel B: 2010–2019 Sample (41,353 Observations) Variable Mean SD Min Max Median Log Labor 9.43 1.66 2.3 15.64 9.37 Log Capital Stock 16.06 2.43 -.7 24.63 16.07 Log Materials 7.71 3.33 -4.61 22.07 7.34 No. of varieties 1.21 .65 1 10 1 Multiproduct .14 .34 0 1 0
Panel C: 2000–2019 Sample (71,079 Observations) Variable Mean SD Min Max Median Log Labor 9.25 1.6 2.3 15.64 9.06 Log Capital Stock 15.75 2.4 -.7 24.32 15.65 Log Materials 7.44 3.35 -4.61 22.99 7.04 No. of varieties 1.36 1.01 1 28 1 Multiproduct .19 .39 0 1 0
Notes: This table presents summary statistics across sample periods, replicating Table 2 from Orr (2022). The capital stock series is constructed using the perpetual inventory method. Consequently, maximum log capital stock values vary across sample periods due to differences in firm entry timing and subsequent accumulation paths. For instance, the firm with the largest capital stock in the 2010–2019 sample (24.63) enters the panel in 2010, whereas in the full 2000–2019 sample, this firm’s earlier entry in 2008 generates a different accumulation trajectory.
Figure A4: Product Counts: ASI Entries versus Number of Product Codes
(a) 2000–2007 Original Sample
(b) 2010–2019 Sample
(c) 2000–2019 Sample
Notes: This figure reports plant-product counts by sample period, where product counts represent either (1) unique entries in the ASI database, permitting multiple entries per plant-product code, or (2) unique product codes within multi-product plants. The category "10" includes plants reporting ten or more varieties.
Figure A5: Sensitivity of First Stage Demand Estimates to Instrument Construction Threshold
Sample Period: 2010-2019
(a) First Stage Coefficient of $p^{j}_{it}$ on $Z^{g(j)}_{t}$
(b) First Stage Coefficient of $rs^{j|g(j)}_{it}$ on $Z^{g(j)}_{t}$
(c) First Stage Coefficient of $p^{j}_{it}$ on $Z^{-jg}_{it}$
(d) First Stage Coefficient of $rs^{j|g(j)}_{it}$ on $Z^{-jg}_{it}$
Notes: This figure displays changes in first-stage estimates reported in Table 3 of Orr (2022) as the instrument construction threshold varies in 0.01 increments. The y-axis presents distinct measures across panels: Panel (a) shows the estimated first-stage coefficient of $p^{j}_{it}$ on $Z^{g(j)}_{t}$, Panel (b) displays the estimated first-stage coefficient of $rs^{j|g(j)}_{it}$ on $Z^{g(j)}_{t}$, Panel (c) reports the estimated first-stage coefficient of $p^{j}_{it}$ on $Z^{-jg}_{it}$, and Panel (d) reports the estimated first-stage coefficient of $rs^{j|g(j)}_{it}$ on $Z^{-jg}_{it}$. All panels include 90% confidence intervals constructed using robust standard errors clustered by plant and product. A vertical dashed line marks the threshold employed in Orr (2022) (0.3). The calculations utilize the 2010–2019 Indian Annual Survey of Industries (ASI) dataset from the Ministry of Statistics and Programme Implementation (MOSPI).
Figure A6: Sensitivity of First Stage Demand Estimates to Instrument Construction Threshold
Sample Period: 2000-2019
(a) First Stage Coefficient of $p^{j}_{it}$ on $Z^{g(j)}_{t}$
(b) First Stage Coefficient of $rs^{j|g(j)}_{it}$ on $Z^{g(j)}_{t}$
(c) First Stage Coefficient of $p^{j}_{it}$ on $Z^{-jg}_{it}$
(d) First Stage Coefficient of $rs^{j|g(j)}_{it}$ on $Z^{-jg}_{it}$
Notes: This figure displays changes in first-stage estimates reported in Table 3 of Orr (2022) as the instrument construction threshold varies in 0.01 increments. The y-axis presents distinct measures across panels: Panel (a) shows the estimated first-stage coefficient of $p^{j}_{it}$ on $Z^{g(j)}_{t}$, Panel (b) displays the estimated first-stage coefficient of $rs^{j|g(j)}_{it}$ on $Z^{g(j)}_{t}$, Panel (c) reports the estimated first-stage coefficient of $p^{j}_{it}$ on $Z^{-jg}_{it}$, and Panel (d) reports the estimated first-stage coefficient of $rs^{j|g(j)}_{it}$ on $Z^{-jg}_{it}$. All panels include 90% confidence intervals constructed using robust standard errors clustered by plant and product. A vertical dashed line marks the threshold employed in Orr (2022) (0.3). The calculations utilize the 2000–2019 Indian Annual Survey of Industries (ASI) dataset from the Ministry of Statistics and Programme Implementation (MOSPI).
Table A5: Robust Replication of Table 4 of Orr (2022): Cobb-Douglas Production Function Estimates from 2010–2019
OLS GMM
(1) (2) (3)
$\beta_{L}$ 0.551 0.691 0.691 0.629
(0.025) (0.052) (0.050) (0.179)
$\beta_{K}$ 0.198 0.080 0.079 0.069
(0.016) (0.021) (0.022) (0.034)
$\beta_{M}$ 0.179 0.291 0.293 0.352
(0.009) (0.032) (0.032) (0.166)
$\rho^{74}$ 0.634 0.634 0.629
(0.025) (0.027) (0.029)
$\rho^{75}$ 0.646 0.646 0.641
(0.022) (0.022) (0.026)
$\rho^{76}$ 0.559 0.559 0.556
(0.036) (0.036) (0.041)
$\rho^{77}$ 0.647 0.648 0.644
(0.019) (0.019) (0.022)
$\rho^{78}$ 0.709 0.707 0.698
(0.053) (0.049) (0.066)
RTS 0.929 1.062 1.063 1.050
(0.016) (0.026) (0.024) (0.044)
Ins.
$(Z_{t}^{g(j)},Z_{t-1}^{g(j)})$ $\checkmark$ $\checkmark$
$m_{it-1}$ $\checkmark$ $\checkmark$
Observations 16,714 16,714 16,714 16,714
  • Notes: Calculations are based on the 2010–2019 Indian ASI dataset, provided by the Ministry of Statistics and Programme Implementation (MOSPI). Each observation represents a plant and a 5-digit ASICC variety. The sample is restricted to producers within the machinery, equipment, and parts industry. The dependent variable is the log quantity of product $j$ produced by plant $i$. Plant-level block bootstrapped standard errors are presented in parentheses.
Table A6: Robust Replication of Table 4 of Orr (2022): Cobb-Douglas Production Function Estimates from 2000–2019
OLS GMM
(1) (2) (3)
$\beta_{L}$ 0.528 0.506 0.508 0.598
(0.026) (0.000) (0.001) (0.004)
$\beta_{K}$ 0.196 0.060 0.060 0.077
(0.017) (0.000) (0.000) (0.001)
$\beta_{M}$ 0.205 0.466 0.465 0.368
(0.009) (0.001) (0.001) (0.005)
$\rho^{74}$ 0.695 0.695 0.702
(0.000) (0.000) (0.000)
$\rho^{75}$ 0.659 0.659 0.665
(0.000) (0.000) (0.001)
$\rho^{76}$ 0.594 0.594 0.597
(0.000) (0.000) (0.000)
$\rho^{77}$ 0.617 0.617 0.624
(0.000) (0.000) (0.000)
$\rho^{78}$ 0.616 0.616 0.640
(0.000) (0.000) (0.001)
RTS 0.930 1.033 1.033 1.043
(0.016) (0.000) (0.000) (0.000)
Ins.
$(Z_{t}^{g(j)},Z_{t-1}^{g(j)})$ $\checkmark$ $\checkmark$
$m_{it-1}$ $\checkmark$ $\checkmark$
Observations 23,185 23,185 23,185 23,185
  • Notes: Calculations are based on the 2000–2019 Indian ASI dataset, provided by the Ministry of Statistics and Programme Implementation (MOSPI). Each observation represents a plant and a 5-digit ASICC variety. The sample is restricted to producers within the machinery, equipment, and parts industry. The dependent variable is the log quantity of product $j$ produced by plant $i$. Plant-level block bootstrapped standard errors are presented in parentheses.
Figure A7: Negative Correlation between TFPQ and Product Appeal
Sample Period: 2010–2019
Notes: This figure replicates Figure 1 from Orr (2022). Appeal refers to $\eta^{j}_{it}$, while TFPQ refers to $\omega^{j}_{it}$. The calculations employ the 2010–2019 Indian Annual Survey of Industries (ASI) dataset from the Ministry of Statistics and Programme Implementation (MOSPI). Both variables are standardized to have zero mean and unit variance within each product code. Panel (1) includes all products, while Panel (2) restricts to single-product plants. Plant-level block bootstrapped standard errors appear in parentheses (1,000 replications).
Figure A8: Negative Correlation between TFPQ and Product Appeal
Sample Period: 2000–2019
Notes: This figure replicates Figure 1 from Orr (2022). Appeal refers to $\eta^{j}_{it}$, while TFPQ refers to $\omega^{j}_{it}$. The calculations employ the 2000–2019 Indian Annual Survey of Industries (ASI) dataset from the Ministry of Statistics and Programme Implementation (MOSPI). Both variables are standardized to have zero mean and unit variance within each product code. Panel (1) includes all products, while Panel (2) restricts to single-product plants. Plant-level block bootstrapped standard errors appear in parentheses (1,000 replications).
Figure A9: Plant-Level Revenue Efficiency Growth After Removing Marginal Varieties
Sample Period: 2010–2019
Notes: This figure replicates Figure 2 from Orr (2022). The figure displays the implied log revenue efficiency growth from removing a plant’s lowest efficiency variety according to core-competence technology. The calculations employ the 2010–2019 Indian Annual Survey of Industries (ASI) dataset from the Ministry of Statistics and Programme Implementation (MOSPI). The y-axis denotes the change in plant-level log revenue efficiency after dropping a plant’s worst product.
Figure A10: Plant-Level Revenue Efficiency Growth After Removing Marginal Varieties
Sample Period: 2000–2019
Notes: This figure replicates Figure 2 from Orr (2022). The figure displays the implied log revenue efficiency growth from removing a plant’s lowest efficiency variety according to core-competence technology. The calculations employ the 2000–2019 Indian Annual Survey of Industries (ASI) dataset from the Ministry of Statistics and Programme Implementation (MOSPI). The y-axis denotes the change in plant-level log revenue efficiency after dropping a plant’s worst product.
Table A7: Product-Dropping Probits (4,907 Observations)
Sample Period: 2010–2019
(1) (2) (3) (4) (5) (6) (7) (8)
$E^{j}_{it}$ -0.0922
$\bar{P}^{g}_{t}$ -0.1052 -0.1028
(0.0169) (0.0142)
$\bar{A}^{j}_{it}$ -0.0379 -0.0312
(0.0162) (0.0159)
$\bar{C}^{j}_{it}$ -0.0228 -0.0289
(0.0139) (0.0132)
$\hat{P}^{g}_{t}$ -0.0595 -0.1072
(0.0165) (0.0271)
$\hat{A}^{j}_{it}$ -0.0754 -0.1379
(0.0232) (0.0317)
$\hat{C}^{j}_{it}$ 0.0132 -0.0623
(0.0106) (0.0253)
  • Notes: This table replicates Table 9 from Orr (2022), presenting marginal effects from a probit regression where the dependent variable indicates whether a plant discontinues production of a specific product in the subsequent period. The analysis uses plant–product observations from India’s Annual Survey of Industries (ASI) dataset spanning 2010–2019, obtained from the Ministry of Statistics and Programme Implementation (MOSPI). The sample includes all producers within the machinery, equipment, and parts industry, with each observation representing a unique plant–product combination at the 5-digit ASICC classification level. Standard errors (in parentheses) are clustered at the plant level using block bootstrap with 1,001 replications.
Table A8: Product-Dropping Probits (15,428 Observations)
Sample Period: 2000–2019
(1) (2) (3) (4) (5) (6) (7) (8)
$E^{j}_{it}$ -0.0796
$\bar{P}^{g}_{t}$ -0.0707 -0.0637
(0.0239) (0.0232)
$\bar{A}^{j}_{it}$ -0.0468 -0.0394
(0.0124) (0.0098)
$\bar{C}^{j}_{it}$ -0.0001 -0.0043
(0.0078) (0.0100)
$\hat{P}^{g}_{t}$ -0.0644 -0.1094
(0.0165) (0.0181)
$\hat{A}^{j}_{it}$ -0.0798 -0.1269
(0.0170) (0.0159)
$\hat{C}^{j}_{it}$ 0.0172 -0.0542
(0.0069) (0.0251)
  • Notes: This table replicates Table 9 from Orr (2022), presenting marginal effects from a probit regression where the dependent variable indicates whether a plant discontinues production of a specific product in the subsequent period. The analysis uses plant–product observations from India’s Annual Survey of Industries (ASI) dataset spanning 2000-2019, obtained from the Ministry of Statistics and Programme Implementation (MOSPI). The sample includes all producers within the machinery, equipment, and parts industry, with each observation representing a unique plant–product combination at the 5-digit ASICC classification level. Standard errors (in parentheses) are clustered at the plant level using block bootstrap with 1,001 replications.